OpenStudy (logan13):

A square garden has an area of 103 square feet. To the nearest tenth, what is the length of one side of the garden? A. 51.5 feet B. 25.8 feet C. 10.1 feet D. 7.2 feet

OpenStudy (sand-lock53):

What is the square root of 103?

OpenStudy (logan13):

10.14

OpenStudy (sand-lock53):

good! now ignore the 4 in 10.14, and you have the right answer.

OpenStudy (phi):

you have to know the area of a square is the length of a side times itself: s^2 = A in this case s^2 = 103 s= sqr(103) we know (right?) 10*10 is 100, so we would expect s to be close to 10

OpenStudy (logan13):

V

OpenStudy (logan13):

*C

OpenStudy (phi):

yes. as a check, you can figure out the area if the side is 10.1 10.1 * 10.1 = 102.01 and that is close to 103 (it's not exact because we rounded to only the nearest 10th)